You are responsible for providing bulk pharmaceutical intermediates to a firm working on a COVID-19 vaccine . The customer specifies that these materials be packaged in one hundred gram portions, with a tolerance of one half gram (100.0 +/- 0.5 g). You sample your output production and find that it has a mean of 100.3 and a standard deviation of 0.12g.

What is your Cp?

What is your Cpk?

Enter your answer rounded to two decimal places
Answer #1

Cp= USL - LSL/ 6 x standard deviation

= 100.5 - 99.5 / 6 x 0.12

= 1 / 0.72

= 1.39

Therefore Cp is 1.39

Cpl = Process mean - LSL / 3 x standard deviation

= 100.3 - 99.5 / 3 x 0.12

= 0.8 / 0.36

= 2.22

Cpu = USL - Process mean/ 3 x standard deviation

= 100.5 - 100.3 / 3 x 0.12

= 0.2 / 0.36

= 0.56

Cpk = min ( 2.22 , 0.56 )